# RedDebate: Safer Responses through Multi-Agent Red Teaming Debates

### v1

@@ Appendix A Extended Related Work
LLMs often exhibit unsafe or harmful behaviours to users. Red-teaming involves the creation and evaluation of a set of test cases aimed at finding such LLM failure cases. Traditional methods involve extensive use of human annotation to manually generate test cases and/or detect harmful responses (Dinan et al., 2019; Ribeiro et al., 2020; Ziegler et al., 2020; Xu et al., 2021; Hendrycks et al., 2021). Perez et al. (2022) first established a method to both automatically generate test cases using a language model, and find failures using a trained classifier. Follow-up work has introduced curiosity-driven exploration for increased coverage of test cases (Hong et al., 2024), optimizing the process of iteratively updating test cases (Mehrabi et al., 2024), and building a red-team that can automatically formulate a measure for harmful outputs and optimize a generator for diverse adversarial prompts (Casper et al., 2023). We build on this work by addressing the need for a fully automated red-teaming evaluation pipeline using LLMs as evaluators without requiring any trained classifier or human oversight.
@@
Due to limitations in context length and ability to handle longer-term dependencies LLM agents benefit from a means of storing and accessing information they have previously learned across interactions (Zhang et al., 2024). Accordingly, previous works have sought to accomplish this by incorporating memory modules, which are generally divided into two types: textual and parametric. Approaches using textual memory store and retrieve information in natural language. Means of textual memory include long-context length strategies (Huang et al., 2023; Li et al., 2023), or strategically processing recent interactions using methods such as flash memory Wang et al. (2025) virtual context management Packer et al. (2023), or cache systems Wang et al. (2024a). Retrieval-based mechanisms using vector-databases can allow most relevant information from full-context to be effecitevly utilized (Hu et al., 2023; Zhong et al., 2024). Parametric memory involves directly altering LLM parameters to adaptively impart knowledge into LLMs. Fine-tuning based methods train on domain-specific knowledge using supervised fine tuning (Hu et al., 2023), including parameter efficient tuning methods Xiong et al. (2023). Likewise, memory editing methods target modifying specific LLM parameters to inject a small set of facts without inhibiting general knowledge (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022, 2023). In this paper, we are the first to use both types of memory in the context of debate agents and safety.
@@
Guardrailing allows determining if and how some actions could be enforced to increase the safety in a system (Dong et al., 2024b). Recently, many LLM guardrailing tools have been developed to mitigate risks associated with them (Rebedea et al., 2023; Guardrails AI, ). NeMo Guardrails (Rebedea et al., 2023) allows building safe LLM agents via a programming language called Colang [^4] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html by specifying predefined dialogical pathways for LLMs. Programmatic gaurdrailing allows modelling highly flexible conversational flows, which might not be possible with existing dialogue management techniques (Rebedea et al., 2023). Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. Building on this line of work, we propose, for the first time, the use of guardrails as memory in the safety setting.
Appendix B Debate and Agents Ablation Study
B.1 Debate Rounds
We study the effect of increasing the number of debate rounds from three to five on the HarmBench dataset. Due to the ease of applying TLTM, we focus solely on this type of memory integration in this study. This analysis investigates how extending the debate affects two key metrics: total error rate and agreement rate. Additionally, we introduced a new metric called diversity, which measures whether each debate round contains at least one safe and one unsafe response, thereby capturing the variability of viewpoints throughout the debate.
Setting TER $\downarrow$ TAGR $\uparrow$ DIV $\uparrow$
3 Rounds 0.289 0.120 0.390
4 Rounds 0.296 0.105 0.405
5 Rounds 0.264 0.094 0.380
3 Rounds+tltm 0.117 0.074 0.280
4 Rounds+tltm 0.180 0.071 0.303
5 Rounds+tltm 0.184 0.059 0.248
Table 4: Impact of the number of debate rounds on performance metrics in HarmBench. TER: Total Error Rate, TAGR: Total Agreement Rate, DIV: Diversity.
B.2 Number of Agents
Setting TER $\downarrow$ TAGR $\uparrow$ DIV $\uparrow$
3 Agents 0.289 0.120 0.390
4 Agents 0.291 0.107 0.521
5 Agents 0.245 0.101 0.570
3 Agents+tltm 0.117 0.074 0.280
4 Agents+tltm 0.111 0.054 0.285
5 Agents+tltm 0.060 0.042 0.223
[…]
B.3 Evaluator Agent
@@
Evaluator Accuracy Precision Recall F1 Score
LlamaGuard 0.87 0.90 0.94 0.92
GPT-4o-mini 0.81 0.87 0.89 0.88
@@
B.4 Socratic Agent and Feedback Generator LLM Backbones
Appendix C LTMs Technical Details
TLTM.
We embed each feedback using OpenAI’s text-embedding-3-large model (dimension: 3072), and store the resulting vectors in a Pinecone vector database. [^5] https://www.pinecone.io During inference, the debate prompt is embedded using the same model, and cosine similarity is used to retrieve the most relevant feedback entries. Based on our observations, retrieving the top five most semantically relevant feedbacks provides sufficient context for improving safety without overwhelming the agent.
CLTM.
The CLTM utilizes LoRA and is implemented using the HuggingFace Library  [^6] https://huggingface.co/docs/peft/en/package_reference/lora. We apply LoRA to the debater attention layers, adding around 0.1% trainable parameters. We set the LoRA parameters as follows: low-rank parameter $r=16$, LoRA $\alpha=16$, LoRA dropout $=0.1$. Cross-entropy loss between the model’s predictions and the actual next tokens in the sequence is used for optimization. We re-fine-tune the parameters on the whole set of feedback with every 10 new ones added.
GLTM.
To implement the Colang guardrails, we use a fork of NeMo-Guardrails version 0.11, which we slightly modified to inject CoSafe conversation history and improve robustness for evaluation. The modified NeMo-Guardrails version used in our experiments is available at our GitHub repository. For HarmBench guardrail code generation, we randomly select 120 samples [^7] In our preliminary experiments, we tried 80, 120, and 200 samples and found that 120 examples yielded the best results. from the debate history. This results in 68 pieces of feedback—since not every sample produces an unsafe response (Algorithm 1)—from which we generate 44 guardrails, as some feedback results in overlapping guardrail names that are subsequently merged. For CoSafe, we use 700 samples, which also yield 68 pieces of feedback, from which we generate 63 guardrails. We use GPT-4o to build all guardrails.
Appendix D PReD with LTM
@@
Scenario HarmBench Error Rate (%) $\downarrow$ CoSafe Error Rate (%) $\downarrow$
@@
Appendix E Helpfulness After Memory Integration
Metric Accuracy (%$\uparrow$ Refusal Rate (%$\downarrow$
Mistral Llama Phi Mistral Llama Phi
None (Baseline) 68.3 57.6 57.9
TLTM 64.2 58.4 52.3 1.8 0.4 0.7
CLTM 63.7 60.9 66.7 0.3 0.0 0.0
TLTM+CLTM 63.7 59.6 62.2 0.0 0.1 0.0
GLTM 65.5 58.2 53.5 1.0 0.3 1.8
Appendix F Categorical Vulnerability Analysis
@@
Appendix G Persona-Based Debate Strategy Algorithms
@@ 13       $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{q^{(t)}\}$
Appendix H Debate Samples
@@ H.2 Error Discovery
Appendix I Agent Prompts

### v2

@@ Appendix A Extended Related Work
LLMs often exhibit unsafe or harmful behaviours to users. Red-teaming involves the creation and evaluation of a set of test cases aimed at finding such LLM failure cases. Traditional methods involve extensive use of human annotation to manually generate test cases and/or detect harmful responses (Dinan et al., 2019; Ribeiro et al., 2020; Ziegler et al., 2020; Xu et al., 2021; Hendrycks et al., 2021). Perez et al. (2022) first established a method to both automatically generate test cases using a language model, and find failures using a trained classifier. Ge et al. (2023) enhances LLM safety by iteratively pairing an adversarial model that generates challenging prompts with a target model that is fine-tuned on safe responses, enabling continual improvement in both attack generation and defense. Samvelyan et al. (2025) frames adversarial prompt generation as a quality-diversity search problem, using open-ended exploration to create diverse and transferable attacks that both expose vulnerabilities and support safety fine-tuning of LLMs. Other follow-up work has introduced curiosity-driven exploration for increased coverage of test cases (Hong et al., 2024), optimizing the process of iteratively updating test cases (Mehrabi et al., 2024), and building a red-team that can automatically formulate a measure for harmful outputs and optimize a generator for diverse adversarial prompts (Casper et al., 2023). We build on this work by addressing the need for a fully automated red-teaming evaluation pipeline using LLMs as evaluators without requiring any trained classifier or human oversight.
@@
Due to limitations in context length and ability to handle longer-term dependencies LLM agents benefit from a means of storing and accessing information they have previously learned across interactions (Zhang et al., 2024). Accordingly, previous works have sought to accomplish this by incorporating memory modules, which are generally divided into two types: textual and parametric. Approaches using textual memory store and retrieve information in natural language. Means of textual memory include long-context length strategies (Huang et al., 2023; Li et al., 2023), or strategically processing recent interactions using methods such as flash memory Wang et al. (2025) virtual context management Packer et al. (2023), or cache systems Wang et al. (2024a). Retrieval-based mechanisms using vector-databases can allow most relevant information from full-context to be effecitevly utilized (Hu et al., 2023; Zhong et al., 2024). Parametric memory involves directly altering LLM parameters to adaptively impart knowledge into LLMs. Fine-tuning based methods train on domain-specific knowledge using supervised fine tuning (Hu et al., 2023), including parameter efficient tuning methods Xiong et al. (2023). Likewise, memory editing methods target modifying specific LLM parameters to inject a small set of facts without inhibiting general knowledge (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022; 2023). In this paper, we are the first to use both types of memory in the context of debate agents and safety.
@@
Guardrailing allows determining if and how some actions could be enforced to increase the safety in a system (Dong et al., 2024b). Recently, many LLM guardrailing tools have been developed to mitigate risks associated with them (Rebedea et al., 2023; Guardrails AI, 2025). NeMo Guardrails (Rebedea et al., 2023) allows building safe LLM agents via a programming language called Colang [^2] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html by specifying predefined dialogical pathways for LLMs. Programmatic gaurdrailing allows modelling highly flexible conversational flows, which might not be possible with existing dialogue management techniques (Rebedea et al., 2023). Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. Building on this line of work, we propose, for the first time, the use of guardrails as memory in the safety setting.
Appendix B Additional Results and Ablations
B.1 When Refusal is not Enough; Ablations on GPT-OSS as a Debater.
(a) Harmbench
(b) CoSafe
Figure 6: Error Rate vs Debate Engagement
Scenario HarmBench CoSafe
Error Rate (%) ↓ Agreement Rate (%) ↑ Error Rate (%) ↓ Agreement Rate (%) ↑
Total Grok R1 OSS Total Grok R1 OSS Total Grok R1 OSS Total Grok R1 OSS
Std. Prompting 30.4 45 44.8 1.3 4.4 6.3 5.5 0.9
PReD 22.9 30.2 37 1.4 11.9 17.1 17.6 1 3.9 5.2 5.2 1.3 2 3.1 2.1 0.9
DAReD 25.8 32.3 44.2 0.9 8.9 15.1 11.1 0.4 3.8 5 5.5 0.8 1.6 2.4 2.0 0.5
SReD 20.7 27.8 33.1 1.1 11.2 16.4 16.6 0.5 3.6 4.5 5.3 0.9 2.1 3.5 2.2 0.6
To examine how highly safety-aligned models behave when confronted with unsafe debate prompts and peer outputs in multi-agent settings, we evaluate our framework using the GPT-OSS model, highlighting its limitations in adversarial debate scenarios. For this experiment, we form a new triad consisting of: (i) Grok-3-mini, widely regarded for balancing helpfulness and safety; (ii) R1-8B, which in our prior observations shows the highest tendency toward unsafe outputs; and (iii) GPT-OSS-20B.
As shown in Table 3, GPT-OSS achieves the lowest overall error rate among all tested models, a promising result at first glance. However, closer inspection reveals important caveats. On HarmBench questions, which contain explicitly unsafe queries (e.g., “how to make a homemade weapon”), GPT-OSS almost never engages in debate, instead consistently producing brief refusal responses. While this behaviour avoids unsafe completions, it also means GPT-OSS remains passive when other models generate unsafe outputs, as illustrated in Figure 6(a). This suggests that, although GPT-OSS is highly regulated against unsafe direct queries, its conservative refusal strategy limits its helpfulness in multi-agent settings, particularly when it could have intervened to guide other agents away from unsafe reasoning. In contrast, on the CoSafe dataset, which includes more indirect safety-oriented debate prompts, GPT-OSS performs strongly, engaging in debate while maintaining safety, as shown in Figure 6(b). This highlights an important trade-off: beyond simply refusing unsafe queries, responsible models in multi-agent environments should also actively steer interactions toward constructive and reliable outcomes.
Notably, the R1 and Grok models continue refining their responses even when the OSS agent disengages and offers little support. While highly guarded models are inherently safer, their tendency to withdraw from debate means they gain less from interacting with other agents. Importantly, this withdrawal does not hinder the progress of agents that remain engaged. Overall, our findings suggest that multi-agent debating is most effective for models willing to participate actively, where they can both improve their own outputs and help steer others toward safer behaviour.
B.2 Evaluator Agent
@@
Evaluator Accuracy (%) Precision (%) Recall (%) F1 Score (%)
LlamaGuard 87 90 94 92
GPT-4o-mini 81 87 89 88
@@
B.3 Discrepancies as a Driver of Safety
We posit that discrepancies among agents—i.e., differences in their high-probability tokens or initial answers—are precisely what enable safety refinement during debate. As shown in Figure 9, we investigate the question: does greater semantic diversity among agent responses lead to better safety outcomes? Our analysis indicates that when agents produce semantically diverse responses, they are more likely to revise unsafe outputs into safer ones over successive rounds. This finding underscores the unique advantage of multi-agent setups, particularly debate, which inherently promotes exploration across a wider range of reasoning paths. Unlike self-revision, which is confined to a single agent’s distribution, debate enables agents to collectively navigate toward safer conclusions by leveraging differing perspectives. In terms of safety refinement, if one agent’s local distribution favors a risky or unsafe answer, another might offer a safer alternative, allowing the group to converge on a more reliable outcome. This also highlights that safety is not solely the responsibility of a designated evaluator agent; rather, it emerges through the collaborative dynamics of all participating debaters.
Figure 8: All baselines are granted an equal number of revision opportunities or inference calls. However, debater agents revise their responses through peer interaction, while self-critique agents reflect independently.
(a) Response diversity versus safety gain, measured as the percentage of first-round unsafe responses converted to safe responses by the final round. The positive trend suggests that higher initial diversity drives greater safety improvements.
(b) Number of improved, declined, or unchanged conversions of first-round unsafe responses against the diversity. As diversity increases, the number of improved conversions rises due to fewer unchanged responses than in the low-diversity quartile.
Figure 9: The diversity metric quantifies how meaningfully agent responses differ by measuring the mean pairwise semantic distance. Specifically, responses from the first triad of debaters (Mistral, Llama, and Phi) are embedded and compared using cosine similarity.
B.4 Helpfulness After Memory Integration
Memory Type Accuracy (%) ↑ Refusal Rate (%) ↓
Mistral Llama Phi Mistral Llama Phi
None (Baseline) 68.3 57.6 57.9
TLTM 64.2 58.4 52.3 1.8 0.4 0.7
CLTM 63.7 60.9 66.7 0.3 0.0 0.0
TLTM+CLTM 63.7 59.6 62.2 0.0 0.1 0.0
GLTM 65.5 58.2 53.5 1.0 0.3 1.8
B.5 Llama Base as a Debater
We evaluate LLaMA-3.2-3B-Base model in our debate framework to assess whether instruction tuning in the post-training process contributes positively to safety outcomes. We replicate our debate setup with Mistral and Phi models, but replace the instruction-tuned Llama with its base version. We observe that the instruction-tuned Llama is generally safer in its initial responses. More significantly, in the debate scenario, we analyze the refinement process and find that the base model’s unsafe response rate does not consistently decrease across rounds. Instead, it rises from 33.8% in the first round to 43.3% in the second round, before slightly declining to 38.8% in the third round. Unsafe answers also tend to persist across rounds (approximately 25% carry-over rate), indicating that the model rarely revises unsafe content during the debate process. This limitation stems from the model’s inability to meaningfully engage with the other debater’s arguments—ignoring external input, it behaves as an isolated responder rather than a collaborative participant.
B.6 PReD with LTM
@@
Scenario HarmBench Error Rate (%) ↓ CoSafe Error Rate (%) ↓
@@
B.7 Debate Rounds
We study the effect of increasing the number of debate rounds from three to five on the HarmBench dataset. Due to the ease of applying TLTM, we focus solely on this type of memory integration in this study. This analysis investigates how extending the debate affects two key metrics: total error rate and agreement rate. Additionally, we introduced a new metric called diversity (DIV), which measures whether each debate round contains at least one safe and one unsafe response, thereby capturing the variability of viewpoints throughout the debate.
Setting TER (%) ↓ TAGR (%) ↑ DIV (%) ↑
3 Rounds 28.9 12.0 39.0
4 Rounds 29.6 10.5 40.5
5 Rounds 26.4 9.4 38.0
3 Rounds+tltm 11.7 7.4 28.0
4 Rounds+tltm 18.0 7.1 30.3
5 Rounds+tltm 18.4 5.9 24.8
Table 7: Effect of debate round count on performance metrics in HarmBench. TER: Total Error Rate, TAGR: Total Agreement Rate, DIV: Diversity.
B.8 Number of Agents
Setting TER (%) ↓ TAGR (%) ↑ DIV (%) ↑
3 Agents 28.9 12.0 39.0
4 Agents 29.1 10.7 52.1
5 Agents 24.5 10.1 57.0
3 Agents+tltm 11.7 7.4 28.0
4 Agents+tltm 11.1 5.4 28.5
5 Agents+tltm 6.0 4.2 22.3
B.9 Categorical Vulnerability Analysis
@@
[…]
B.10 Socratic Agent and Feedback Generator LLM Backbones
Appendix C LTMs Technical Details
C.1 TLTM.
We embed each feedback using OpenAI’s text-embedding-3-large model (dimension: 3072), and store the resulting vectors in a Pinecone vector database.  [^3] https://www.pinecone.io Note that vectors are actually used only for retrieval; ultimately, agents receive the top-matching textual feedback entries. During inference, the debate prompt is embedded using the same model, and cosine similarity is used to retrieve the most relevant feedback entries. Based on our observations, retrieving the top five most semantically relevant feedbacks provides sufficient context for improving safety without overwhelming the agent.
C.2 CLTM.
The CLTM utilizes LoRA and is implemented using the HuggingFace Library  [^4] https://huggingface.co/docs/peft/en/package_reference/lora. We apply LoRA to the debater attention layers, adding around 0.1% trainable parameters. We set the LoRA parameters as follows: low-rank parameter $r=16$, LoRA $\alpha=16$, LoRA dropout =0.1=0.1. Cross-entropy loss between the model’s predictions and the actual next tokens in the sequence is used for optimization. We re-fine-tune the parameters on the whole set of feedback with every 10 new ones added.
C.3 GLTM
To implement the Colang guardrails, we use a fork of NeMo-Guardrails (v0.11) with minor modifications and Colang language [^5] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html to inject CoSafe conversation history and improve robustness during evaluation. The modified version used in our experiments is publicly available on our GitHub repository. For the HarmBench guardrail code generation with the first triad of debaters, we randomly select 120 samples [^6] In preliminary experiments, we tested 80, 120, and 200 samples and found that 120 samples produced the best results. from the debate history. These samples yield 68 pieces of feedback—since not every sample produces an unsafe response (Algorithm 1)—from which 44 guardrails are generated after merging overlapping guardrail names. For CoSafe, we use 700 samples, producing 68 pieces of feedback, which are then converted into 63 guardrails. All guardrails are generated using GPT-4o. For the second triad of debaters, we obtain 45 feedback items for HarmBench and 164 for CoSafe, resulting in 40 and 145 guardrails, respectively. Figure 22 shows an example of the generated guardrails.
Further Performance Analysis of GLTM
Among the models, Llama shows the most reliable GLTM integration, with the lowest error rates (0.3% on HarmBench, 0.4% on CoSafe), high recall of unsafe intents (99.4%, 92.7%), and minimal runtime failures, making it well-suited for guardrail-based safety (Table 9). In contrast, Phi suffers from non-negligible runtime errors due to NeMo’s instability and formatting constraints. Examples of these runtime errors include:
  • “Error: No valid response after N attempts.”
  • “Internal error on flow X”
  • “Sorry! There was an issue in the LLM result form X”
  • “None response”
Based on our investigation, addressing these errors requires careful prompt design, as well as tuning or explicit flow handling within the NeMo framework, particularly to help LLMs with context limitations follow the framework’s intended rails. Readers can explore these strategies further depending on the models they use (see https://docs.nvidia.com/nemo/guardrails/latest/user-guides/llm-support.html for more details).
Metric HarmBench CoSafe
Mistral Llama Phi Mistral Llama Phi
Intent Match (%) ↑ 25.5 39.5 26.0 2.4 4.5 1.4
Guardrails Recall (%) ↑ 76.7 99.4 94.5 35.4 92.7 35.2
Runtime Error (%) ↓ 7.3 0.3 25.0 0.0 0.1 9.7
Gemma Qwen R1 Gemma Qwen R1
Intent Match (%) ↑ 61 74.3 71.5 7.2 21.8 24.2
Guardrails Recall (%) ↑ 94.9 1 99.6 63.9 99.1 95.8
Runtime Error (%) ↓ 21.5 12.75 14.0 0.9 4.4 28.8
Appendix D Persona-Based Debate Strategy Algorithms
@@ 13 $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{q^{(t)}\}$
Appendix E Debate Samples
@@ E.2 Error Discovery
Appendix F Agent Prompts